OpenStudy (anonymous):

Radius OA of circle O has a slope of 2.5. What is the slope of the line tangent to circle O at point A? 0.8 2.5 -0.4 -5

OpenStudy (amoodarya):

|dw:14351729953:dw

OpenStudy (amoodarya):

 \[\frac{HA}{oH}=2.5\] is this ? what you mean ?

OpenStudy (anonymous):

um i guess...it's cause the problem doesn't include a picture

OpenStudy (amoodarya):



OpenStudy (amoodarya):

you can draw an image (it help you to better understanding) slope (oA) , slope(tangent line) =-1 because they are perpendicular

OpenStudy (amoodarya):

now you must guess the answer ! can't you ?

OpenStudy (anonymous):

i dont know! i honestly dont know /:

OpenStudy (amoodarya):

if two line are perpendicular ...with slope \[m_1 ,m_2\] so relation between them is : \[m_1.m_2=-1\\m_1=2.5\\so\\m_2=?\]

OpenStudy (amoodarya):

m2 =?

OpenStudy (anonymous):

-3.5? idk!

OpenStudy (amoodarya):

\[2.5 *m_2=-1\\m_2=\frac{-1}{2.5}=\frac{-1}{\frac{5}{2}}\\=\frac{-2}{5}\]

OpenStudy (anonymous):

-0.4

OpenStudy (amoodarya):

bravo

OpenStudy (anonymous):

yaaay! thank you!

OpenStudy (amoodarya):

understand ?

OpenStudy (anonymous):

yes (:

OpenStudy (amoodarya):

it is not better to change your Id to igetmath ?

OpenStudy (anonymous):

you dont like my id name? lol

OpenStudy (amoodarya):

i live mathematics